# OmniDiff: A Comprehensive Benchmark for Fine-grained Image Difference Captioning

### v1

@@ 3.1 Overview of OmniDiff
We introduce OmniDiff, a high-quality dataset encompassing diverse real-world and synthetic complex scenarios, featuring fine-grained human annotations to ensure comprehensive and accurate labeling. OmniDiff comprises 324 scene pairs spanning diverse indoor and outdoor environments, collected through a combination of on-site photography, web scraping and 3D rendering. The dataset encompasses 12 categories of fine-grained visual variations, including viewpoint, illumination, addition, disappearance, removal, substitution, size, color, orientation, pose, OCR (textual changes), and counting. Each similar image pair encompasses one or more types of variations, with the average length of difference descriptions extending to 60 words to ensure a thorough and detailed representation of the observed changes.
@@
Table 1: The comparison of OmniDiff with existing image difference caption datasets. OmniDiff excels in both its breadth, encompassing a diverse range of real-world scenarios, and its depth, characterized by detailed and fine-grained annotations. The underlined value indicates a relatively high number of image pairs in CLEVR-Change, although these pairs are limited to tabletop environments with constrained object variations. Distractor change refers to non-semantic variations (e.g.,viewpoint shifts and illumination).
@@ 3.2 Dataset Construction Process
The construction process of the OmniDiff consists of two key stages: Change Pair Collection and Difference Caption Collection. The overall workflow architecture is illustrated in Figure 2.
@@
To collect change pairs from real-world scenes, we utilize a combination of on-site photography and web crawling to capture diverse environmental conditions and various types of changes, ensuring comprehensive coverage of real-world scenarios. For on-site photography, we systematically position cameras in diverse indoor and outdoor environments to capture natural scene variations over time, while adjusting camera perspectives to simulate viewpoint changes and enrich the diversity of the collected data. To further enhance scene diversity, we collect online videos to capture environments and change types that are challenging to obtain through on-site photography. We specifically select videos depicting real-life scenarios and manually extract frames with similar scenes but subtle visual differences to construct change pairs. The sources and quality of these videos are rigorously screened to ensure compliance with the dataset’s high-quality standards. Through a combination of on-site photography and web crawling, we collect change pairs from 224 distinct scenes.
@@ 4.2 Multi-Scale Differential Perception
[…]
@@
$\text{Score}^{i}$ $=\sigma\Big{(}\text{MLP}\big{(}\Phi_{\text{Mean}}(\tilde{\mathbf{F}}_{1}^{i}\|\tilde{\mathbf{F}}_{2}^{i}\|\Delta\mathbf{F}^{i})\big{)}\Big{)},$ (7)
@@ 5.1 Datasets
Image Editing Request  [36] consists of 3,939 aligned image pairs from real-life scenarios, accompanied by 5,695 editing instructions. Following the official dataset split protocol [36], we use 3,061 image pairs for training, 383 for validation, and 495 for testing.
@@ 5.2 Implementation Details
Architecture. The model initializes the weights of the visual encoder, projector, and LLM from LLaVA-OneVision-7B [20], while the MDP module weights are newly initialized. The visual encoder utilizes the SigLIP model [49], which consists of 27 transformer layers. Multi-scale features extracted from layers 17, 20, 23, and 26 of the visual encoder serve as inputs to the MDP module, which comprises two stacked transformer layers. The projector comprises a two-layer multilayer perceptron (MLP), and the LLM is built upon the Qwen2 architecture [47].
@@
Table 5: Performance comparison on Image-Editing-Request across all metrics. The best results are highlighted in boldface, while the second-best results are indicated with underlining.
@@ 5.3 Performance Comparison
To validate the effectiveness of M3Diff in complex and dynamic scenarios, we fine-tune state-of-the-art IDC methods (VARD [39], SCORER [41], DIRL [42], CARD [44]) and the MLLM-based approach FINER-MLLM [50] on the OmniDiff dataset. We then conduct a comprehensive evaluation on the test set to compare their performance. As demonstrated in Table 3, M3Diff achieves superior performance across all metrics in both real-world and rendered scenarios. Additionally, we perform zero-shot evaluations of leading MLLMs (GPT-4o [30], LLaVA-OneVision [20], Qwen-2.5-VL [4]) on OmniDiff, demonstrating their limitations in analyzing complex scene differences despite their general multimodal capabilities.
As shown in Tables 4, 5, 7, and 8, M3Diff consistently achieves state-of-the-art performance across various publicly IDC benchmarks, including real-world scenarios such as Spot-the-Diff [15] and IEdit [36], as well as rendered environments like CLEVR-Change [32] and CLEVR-DC [18]. These results demonstrate the model’s effectiveness and strong generalization ability in cross-scenario difference identification.
@@ 5.4 Ablation Study and Analysis
Effects of Multi-Scale Differential Perception Module. Table 6 demonstrates that introducing the MDP module to enhance difference perception capabilities, while preserving the inherent knowledge of LLaVA-OneVision [20], leads to further improvements on both the OmniDiff and IEdit [36] test set. Removing MDP leads to consistent degradation in text quality and semantic coherence, highlighting its role in harmonizing multi-scale feature integration. While partial metric fluctuations occur (i.e., the CIDEr variation on OmniDiff-Real), the full model demonstrates balanced superiority, outperforming ablated versions in 12/13 metrics. This holistic enhancement confirms MDP’s necessity for fine-grained IDC tasks.
@@ 6 Conclusion and Future Outlook
In future work, the OmniDiff dataset can be further expanded to include image sets capturing continuous real-world changes, enabling more comprehensive modeling of dynamic scenarios. Building on this, M3Diff can further be extended to video-based difference captioning, significantly broadening its applicability to dynamic and temporally evolving environments.
@@
In this supplementary material, we provide additional details regarding the OmniDiff dataset and the training specifics of the M3Diff model. Finally, we analyze the qualitative performance of four models on samples from the OmniDiff dataset.
@@ 8.2 Hyperparameters
9 Case Study
In this section, we aim to analyze and compare the difference captions provided by different models (GPT-4o [30], CARD [44], FINER-MLLM [50], M3Diff) for image pairs in our dataset. The focus is on how each model captures the presence and behavior, as well as other changes in the scene.

### v2

@@ 3.1 Overview of OmniDiff
We introduce OmniDiff, a high-quality dataset encompassing diverse real-world and synthetic complex scenarios, featuring fine-grained human annotations to ensure comprehensive and accurate labeling. OmniDiff comprises 324 scenes spanning diverse indoor and outdoor environments, collected through a combination of on-site photography, web scraping and 3D rendering. The dataset encompasses 12 categories of fine-grained visual variations, including viewpoint, illumination, addition, disappearance, removal, substitution, size, color, orientation, pose, OCR (textual changes), and counting. Each similar image pair encompasses one or more types of variations, with the average length of difference descriptions extending to 60 words to ensure a thorough and detailed representation of the observed changes.
@@
Table 1: The comparison of OmniDiff with existing image difference caption datasets. OmniDiff excels in both its breadth, encompassing a diverse range of real-world scenarios, and its depth, characterized by detailed and fine-grained annotations. The underlined value indicates a relatively high number of image pairs in CLEVR-Change [33], although these pairs are limited to tabletop environments with constrained object variations. Distractor change refers to non-semantic variations (e.g.,viewpoint shifts and illumination changes).
@@ 3.2 Dataset Construction Process
The construction process of the OmniDiff consists of two key stages: Change Pair Collection and Difference Caption Collection. Figure 2 shows the overall workflow.
@@
To collect change pairs from real-world scenes, we utilize a combination of on-site photography and web crawling to capture diverse environmental conditions and various types of changes, ensuring comprehensive coverage of real-world scenarios. For on-site photography, we systematically position cameras in diverse indoor and outdoor environments to capture natural scene variations over time, while adjusting camera perspectives to simulate viewpoint changes and enrich the diversity of the collected data. To further enhance scene diversity, we collect online videos to capture environments and change types that are challenging to obtain through on-site photography. We specifically select videos depicting real-life scenarios and manually extract frames with similar scenes but subtle visual differences to construct change pairs. The sources and quality of these videos are rigorously screened to ensure compliance with the dataset’s high-quality standards. Through a combination of on-site photography and web crawling, we collect change pairs from 224 distinct real scenes.
@@ 4.2 Multi-Scale Differential Perception
where $\mathbf{F}_{1}^{i},\mathbf{F}_{2}^{i}$ are image pair features from the $i$-th layer, $\|$ denotes concatenation, and $\sigma$ is sigmoid function. The weight $\mathbf{W}_{m}\in\mathbb{R}^{d\times 2d}$ processes concatenated features through a learnable projection, where $d$ denotes the feature dimension. The resulting attention weights $\boldsymbol{\lambda}_{k}\in[0,1]^{d}$ modulate input features via element-wise multiplication ($\odot$), preserving original dimensions. $\mathbf{W}_{p}\in\mathbb{R}^{d\times 3d}$ projects the concatenated triplet into $\Delta\mathbf{F}^{i}$, a $d$-dimensional vector that explicitly encodes feature variations.
@@
$\text{Score}^{i}$ $=\sigma\Big(\text{MLP}\big(\Phi_{\text{Mean}}(\tilde{\mathbf{F}}_{1}^{i}\|\tilde{\mathbf{F}}_{2}^{i}\|\Delta\mathbf{F}^{i})\big)\Big),\$(7)
@@ 5.1 Datasets
Image-Edit-Request  [37] consists of 3,939 aligned image pairs from real-life scenarios, accompanied by 5,695 editing instructions. Following the official dataset split protocol [37], we use 3,061 image pairs for training, 383 for validation, and 495 for testing.
@@ 5.2 Implementation Details
Architecture. M3Diff initializes the weights of the visual encoder, projector, and LLM from LLaVA-OneVision-7B [22], while the MDP module weights are newly initialized. The visual encoder utilizes the SigLIP model [50], containing 27 transformer layers. Multi-scale features extracted from layers 17, 20, 23, and 26 of the visual encoder serve as input to MDP, which comprises two stacked transformer layers. The projector comprises a two-layer multilayer perceptron, and the LLM is based on the Qwen2 [48].
@@
Table 5: Performance comparison on Image-Edit-Request across all metrics. The best results are highlighted in boldface, while the second-best results are indicated with underlining.
@@ 5.3 Performance Comparison
To validate the effectiveness of M3Diff in complex and dynamic scenarios, we fine-tune state-of-the-art IDC methods (VARD [38], SCORER [42], DIRL [40], CARD [43]) and the MLLM-based approach FINER-MLLM [51] on the OmniDiff dataset. We then conduct a comprehensive evaluation on the test set to compare their performance. As demonstrated in Table 3, M3Diff achieves superior performance across all metrics in both real-world and rendered scenarios. Additionally, we perform zero-shot evaluations of leading MLLMs (GPT-4o [1], LLaVA-OneVision [22], and Qwen-2.5-VL [5]) on OmniDiff, demonstrating their limitations in analyzing complex scene differences despite their general multimodal capabilities.
As shown in Tables 4, 5, 7, and 8, M3Diff consistently achieves SOTA performance across various publicly IDC benchmarks, including real-world scenarios such as Spot-the-Diff [17] and IEdit [37], as well as rendered environments like CLEVR-Change [33] and CLEVR-DC [20]. These results demonstrate the model’s strong generalization ability in cross-scenario difference identification.
@@ 5.4 Ablation Study and Analysis
Effects of Multi-Scale Differential Perception Module. Table 6 demonstrates that introducing the MDP module to enhance difference perception capabilities, while preserving the inherent knowledge of MLLM, leads to further improvements on both the OmniDiff and IEdit [37] test set. Removing MDP leads to consistent degradation in text quality and semantic coherence, highlighting its role in harmonizing multi-scale feature integration. While partial metric fluctuations occur (i.e., the CIDEr variation on OmniDiff-Real), the full model demonstrates balanced superiority, outperforming ablated versions in 12/13 metrics. This holistic enhancement confirms MDP’s necessity for fine-grained IDC tasks.
@@ 6 Conclusion and Future Outlook
In future work, OmniDiff can be expanded to include image sets capturing continuous real-world changes, enabling more comprehensive modeling of dynamic scenarios. Building on this, M3Diff can further be extended to video-based difference captioning, significantly broadening its applicability to temporally evolving environments.
Acknowledgement
This work is jointly supported by National Natural Science Foundation of China (62206022, 62276025, 62476027) and the Fundamental Research Funds for the Central Universities (2253200026).
@@ Acknowledgement
In this supplementary material, we provide additional details regarding the OmniDiff dataset and the training specifics of the M3Diff model. Finally, we further experimentally validate the effectiveness of the OmniDiff dataset and the MDP module, and analyze the qualitative performance of four models on samples from OmniDiff.
@@ 8.2 Hyperparameters
9 Experiments
9.1 Performance Comparison with Identical Data
For a fair comparison with other MLLM-based IDC methods (e.g., FINER-MLLM [51]), we reproduce this model using the same 896k training data as M3Diff. As shown in Table 11, M3Diff demonstrates consistent superiority across all metrics on the OmniDiff and CLEVR-DC benchmarks.
9.2 Extended Analysis of the MDP Module
To further validate the effectiveness of the plug-and-play MDP module, we conduct experiments using two advanced MLLMs, Qwen-2.5-VL-7B-Instruct [5] and InternVL3-8B-Instruct [53], as our base models. Table 12 demonstrates that integrating the MDP module into various backbones consistently enhances the model performance. Additionally, M3Diff with LLaVA-OneVision-7B [22] as the backbone maintains strong performance, potentially due to the increased use of multi-image task data in pre-training.
9.3 Extended Analysis of the OmniDiff Dataset
With exceptional instruction-following and semantic understanding capabilities, LLMs serve as a crucial tool for measuring sentence similarity [11]. Therefore, to comprehensively validate the impact of the OmniDiff dataset on model performance, we design prompts to instruct GPT-4o [1] to evaluate semantic consistency between predictions and annotations, outputting a score ranging from 0 to 100. The prompt is shown in Figure 6. Table 13 shows that the LLM-based evaluation confirms the effectiveness of OmniDiff.
Table 11: Performance comparison with FINER-MLLM [51] under identical training data.
Method OmniDiff-Real OmniDiff-Render CLEVR-DC BLEU-4 METEOR ROUGE-L CIDEr BLEU-4 METEOR ROUGE-L CIDEr BLEU-4 METEOR ROUGE-L CIDEr SPICE FINER-MLLM 9.4 14.5 26.1 20.6 13.8 15.8 30.5 18.0 53.1 34.2 69.5 92.3 17.8 M3Diff (ours) 14.3 18.9 32.9 31.3 15.7 19.9 35.3 28.3 60.6 37.6 73.0 109.4 21.3
Table 12: Ablation study of the MDP module based on Qwen2.5-VL-7B-Instruct [5] and InternVL3-8B-Instruct [53]. † indicates the model is evaluated in a zero-shot setting.
Method OmniDiff-Real OmniDiff-Render Image-Edit-Request BLEU-4 METEOR ROUGE-L CIDEr BLEU-4 METEOR ROUGE-L CIDEr BLEU-4 METEOR ROUGE-L CIDEr SPICE Qwen2.5-VL-7B† 3.8 9.5 19.8 6.2 2.1 6.8 18.3 3.3 16.0 9.8 23.2 32.3 10.8 Qwen2.5-VL-7B-SFT w/o MDP 13.1 16.7 31.8 35.3 14.9 18.7 34.7 25.2 30.0 26.5 58.3 131.2 26.6 Qwen2.5-VL-7B-SFT with MDP 13.8 18.8 32.9 36.7 15.5 19.7 35.6 27.8 31.2 26.8 59.1 132.5 27.7 InternVL3-8B† 2.7 12.2 19.2 4.3 4.0 9.2 19.3 5.1 12.1 9.2 17.1 28.6 9.3 InternVL3-8B-SFT w/o MDP 12.5 17.0 31.3 34.6 14.7 19.1 35.1 26.1 29.4 25.7 57.1 128.1 26.3 InternVL3-8B-SFT with MDP 13.5 18.5 32.5 35.7 15.5 20.1 35.7 28.1 30.8 26.1 58.2 130.5 27.1 M3Diff (ours) 14.3 18.9 32.9 31.3 15.7 19.9 35.3 28.3 33.6 26.5 59.7 136.6 27.5
Table 13: Ablation study of OmniDiff based on LLM evaluation.
Method OmniDiff-Real OmniDiff-Render Image-Edit-Request LLM LLM LLM M3Diff w/o OmniDiff 12.3 15.8 63.8 M3Diff with OmniDiff 37.3 37.0 68.5
Figure 6: The prompt for LLM-based evaluation.
10 Case Study
In this section, we aim to compare the difference captions provided by different models (GPT-4o [1], CARD [43], FINER-MLLM [51], M3Diff) for image pairs in our dataset. The focus is on how each model captures the presence and behavior, as well as other changes in the scene.